Hama is three times as old as david and the sum of their age is 52. determine their age
Answer by lenny460(1073)

Hama is three times as old as david and the sum of their age is 52. Determine their ages ?

Let the age of David be n
Hama's age is 3n
Therefore:
n + 3n = 52
4n = 52
n = 13
3n = 3 * 13 = 39
13 + 39 = 52

Answer:

Hama is 39 years old and

David is 13 years old.








Lennox Obuong
Algebra Tutor
Nairobi, Kenya